HUBBLE SPACE TELESCOPE DAILY REPORT #3070 PERIOD COVERED: 0000Z (UTC) 03/11/02 - 0000Z (UTC) 03/12/02 Daily Status Report as of 071/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Science operations are scheduled to begin at approximately 071/0122Z, just after the period of this report. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 12 Successful: 12 2.3 Operations Notes: All HST subsystems are reported to be in a nominal state. The Science Instruments are all out of safe and awaiting resumption of post-SM3B verification and normal science. Per operations requests, the following actions were taken: A gyro bias update at 070/1612Z. An adjustment of SM3B limits at 070/1702Z. An attitude reference update at 070/1830Z. Gyro #6 was powered off at 070/2204Z. Gyro 6 limits were then adjusted. Using ROP DF-10A, a 486 safeware/table dump was made at 070/1715Z. SSR-1 uncorrectable EDAC errors were cleared at 070/1934Z per ROP SR-9A. An engineering status buffer limit (SESBSLD=1) was modified at 071/0221Z, using ROP DF-18A. At 071/025129Z, during a guide star acquisition, QSFGENB0 (Saved FGS Error Null Bias 0) and QFGENB0 (FGS error null Bias 0) flagged out-of-limits low. HSTAR 8548 was written. As documented in HSTAR 8549, the NICMOS Cooling System CPL radiator B temperature (MNCRADBT) flagged out-of-limits low with a value of -68.33 , the NICMOS Cooling System CPL Htex2 temperature (MNCHX1T) flagged low at -63.33, and the NICMOS Cooling System CPL Radiator A temperature (MNCRADAT) flagged low at -68.165. All occurrences were from 071/051558Z to 071/052748Z. During the WF/PC-2 decontamination, the limits were managed at 071/0803Z per an operations note. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Commencement of Servicing Mission Orbital Verification (SMOV) activities and the eventual continuation of normal science observations and calibrations. /CAW